What are the values of the minimum number of resistors needed to give the range 1 ohm's to 10 ohm's in integer steps? Is your set of values unique

Answer is 1,2,3,4 (ohms) or 1,2,4,5 (ohms)......How?
Answer #1

Concept= question is asking u to take minimum number of integral value resistors(in series) so as to get all integral value of resistor from 1 to 10(integral again). There is no specific method to solve. Hit n trial would be the best to proceed.

For ex. Let's see the first set. 1,2,3,4

A)in this set we already have resistors from 1 to 4. We have to work for 5 to 10.

5 =1+4, 6=4+2, 7=4+3, 8=4+3+1, 9=4+3+2, 10=4+3+2+1

Note:-All resistors mentioned above are connected in series to get the desired result.

B) similarly we can work to explain the next set

1245

In this set we need 3 and 6 to 10.

3=2+1, 6=5+1, 7=5+2, 8=5+2+1, 9=5+4, 10=5+4+1

Let me know if it was helpful. Thank you!